NAOJ GW Elog Logbook 3.2


-200 lens for reflection telescope is on OPO transmission path and it changes mode matching of OPO transmission. So this configuration is not feasible.

I found a new solution, better than the previous one, considering a larger database of lenses.
The robustness is good, moving one lens in a range of 1cm, the total mismatch is lower than 20%.
-200 lens for reflection telescope is on OPO transmission path and it changes mode matching of OPO transmission. So this configuration is not feasible.

I have found (between BS adn NM1 chamber) a rack with 5 more boards for the galvo control. See attached picture.
Maybe some of them are the "new version" Yuefan was talking about?
On two of them there is also a label specifing if the QPD has big or small range.

After solving the DGS issue with the filter loading, I could test the simulink model on the control of YAW of PR.
The mechanical TF and the closed loop TF are shown in pic 1 and 2. The comparison between the open and closed loop spectrum is shown in pic 3. The control seems to work fine.
UGF is at ~ 6 Hz and phase margin is ~ 50 deg. A first, rather basic version of medm screen developped for the control is shown in pic 4.
Error and correction signals are currenty in counts and needs to be calibrated.

Today we used the sound spectrum analyzer characterized the sound environment. We found a clear frequency from our cleanroom fan. It is 200Hz.

Yuhang, Pierre, Aritomi
In the beginning, we tried to check the alignment of homodyne. But we cannot see a meaningful signal from AMC transmission.
We found the PZT of AMC was broken. I guess it is related with the strong force we(mainly it's me) enforced on this connector or its wire. I am sorry that I made very ugly soldering.
Anyway, we repaired it and we heard the sound of PZT by sending a 4kHz signal.
So in the future, let's be kind for our wires!

I report the robustness of the injection and reflection telescopes described in entry #1296.
The two telescopes can be consiedered robust enough (till 20% of mismath) only in a range of +/- 2 mm.

[T.Yamamoto, Y.Fujii, Eleonora]
Here the report from Yamamoto-san about today's work:
- Real-time models was not able to read filter files.
Models detected the modification of filter files.
But “COEFF LOAD” button did not work well.
- We unified the RCG version as v3.1.1.
At first master and slave model run as v2.8.8 and v3.1.1, respectively.
But the problem was not solved by unifying the version.
- System clock of STDA on BIOS was fixed.
System clock should be set as UTC. But it was set as JST.
So system time showed 9 hours future and date of file modification was wrong.
We fixed the time-stamp of filter file, but problem was not solved.
- Filter files re-generated after fixing system clock.
We moved filter files and re-generated them by rebuilding models.
But the problem was not solved.
- The version of real-time system returned back form v3.1.1 to v2.8.8.
The problem was solved by using v2.8.8.
#####################
Thanks a lot to Yamanoto-san and Fujii-san for the precious help and for all the time spent!

I made new simulations taking into account also the telescope for the beam reflected from the cavity into the homodyne.
Injection telescope:
focal lenght 1 = -101.65 mm
focal lenght 2= 204.6 mm
Reflection telescope:
focal lenght 1= 203.3 mm
focal lenght 2 = -101.65 mm
In fig 3 you can find the scheme of the two telescopes on the bench.
Next step:
- Test the two telescopes

I just measured there are 3uW of p-pol is going also into homodyne.

1-Setting of switches on the front panel:
* The differentiator shall be disabled on the front panel in setting the switch on "OFF".
* The switch INV/NON INV on the front panel, shall be set on INV.
2-Setting of the 8 straps on the board:
Low-pass filter, Notch filter 1 and notch filter 2 are activated on the board in setting strap on connectors P7, P8 and P9 (3 pins) between pins 1 and 2
* The transmission signal is positive with a peak at 1.26V.
It shall be inverted: the strap on connector P4 (3 pins) is set between pin 2 and 3.
The threshold level must normally be tuned to a negative level of 600mV (THRESHOLD OUT).
We had also to increase the sample-hold capacitor (C89). See below.
* Strap is set on connector P11 (3 pins), between pins 2 and 3, in order to activate the sample-and-hold on the triangular signal, on the locking.
* Strap is set on connector P3 (2 pins) to connect the triangular signal to the output stage.
* Strap is set on connector P2 (3 pins), between pins 1 and 2, for test purpose.
To check low-pass filter, notch 1 and notch 2 filters (in scan mode) between TEST IN and TEST OUT. For this test the differentiator, shall be set on "ON" (not intuitive but important). After this test, the differentiator shall be disabled the front panel in setting the switch on "OFF".
* Strap is set on connector P1 (2 pins), in order to be able to tune the offset.
3-Modification of components:
* Integrator 1/f: corner frequency changed to 3.3 kHz
Capacitor CMS 1206: C38 = 22nF
* Integrator 1/f2: corner frequency changed to 220 Hz
Capacitor CMS 1206: C26 = C33 = 330nF
* Low-pass filter: cut-off frequency changed to 3.3 kHz
Capacitor CMS 0805 : C45 = 2.2nF (0805)
Resistor CMS 1206 : R59 = 22k
* Notch filter 1: notch frequency changed to 10.5 kHz / quality factor changed to 6 (measured)
[Capacitor CMS 0805 1% : C49 ; C50 ; C51 ; C53 = unchanged (560 pF)]
Resistor CMS 1206 : R65 ; R66 ; R67 ; R68 = 27k
Resistor CMS 1206 : R73 = 820
* Notch filter 2: notch frequency changed to 14.2 kHz / quality factor changed to 6 (measured)
[Capacitor CMS 0805 1% : C60 ; C61 ; C62 ; C63 = unchanged (560 pF)]
Resistor CMS 1206 : R79 ; R80 ; R81 ; R82 = 20k
Resistor CMS 1206: R89 = 820
* Gain adjustment (G): Gmin = 0.125 / Gmax = 05 / Gtyp = 1
No modification.
* Modification of he sample-hold capacitance on the triangular signal:
Capacitor 1206 of 4.7microFarad added on the 1microFarad capacitor (C89).

1-Setting of switches on the front panel:
* The differentiator shall be disabled on the front panel in setting the switch on "OFF".
* The switch INV/NON INV on the front panel, shall be set on INV.
2-Setting of the 8 straps on the board:
Low-pass filter, Notch filter 1 and notch filter 2 are activated on the board in setting strap on connectors P7, P8 and P9 (3 pins) between pins 1 and 2
* The reflection signal is negative with a base at about -600mV (after the change of the trans-impedance resistor of the photodetector from 51 Ohm to 510 Ohm ).
It shall be inverted: the strap on connector P4 (3 pins) is set between pin 2 and 3.
The threshold level must normally be tuned to a positive level of +300mV (THRESHOLD OUT).
In practice, we had best result with a threshold level of about +500mV.
We had also to increase the sample-hold capacitor (C89). See below.
* Strap is set on connector P11 (3 pins), between pins 2 and 3, in order to activate the sample-and-hold on the triangular signal, on the locking.
* Strap is set on connector P3 (2 pins) to connect the triangular signal to the output stage.
* Strap is set on connector P2 (3 pins), between pins 1 and 2, for test purpose.
To check low-pass filter, notch 1 and notch 2 filters (in scan mode) between TEST IN and TEST OUT. For this test the differentiator, shall be set on "ON" (not intuitive but important). After this test, the differentiator shall be disabled the front panel in setting the switch on "OFF".
* Strap is set on connector P1 (2 pins), in order to be able to tune the offset.
3-Modification of components:
* Integrator 1/f: corner frequency changed to 154 Hz
Capacitor CMS 1206: C38 = 470nF
* Integrator 1/f2: corner frequency changed to 220 Hz
Capacitor CMS 1206: C26 = C33 = 330nF
* Low-pass filter: cut-off frequency changed to 154 Hz
Capacitor CMS 0805 : C45 = 2.2nF (0805) + 270nF (through capacitor)
Resistor CMS 1206 : R59 = 3.9k
* Notch filter 1: notch frequency changed to 11.85 kHz / quality factor changed to 3 (measured)
[Capacitor CMS 0805 1% : C49 ; C50 ; C51 ; C53 = unchanged (560 pF)]
Resistor CMS 1206 : R65 ; R66 ; R67 ; R68 = 24k
Resistor CMS 1206 : R73 = 2.7k
* Notch filter 2: notch frequency changed to 17.75 kHz / quality factor changed to 0.9 (measured)
[Capacitor CMS 0805 1% : C60 ; C61 ; C62 ; C63 = unchanged (560 pF)]
Resistor CMS 1206 : R79 ; R80 ; R81 ; R82 = 16k
Resistor CMS 1206: R89 = 13k
* Gain adjustment (G): Gmin = 0.0125 / Gmax = 0.5 / Gtyp = 0.1ÂÂ
Resistor CMS 1206 : R33 = 2.2k
Resistor CMS 1206 : R5 = 10k
Resistor CMS 1206 : R7 = 10k
* Modification of threshold circuit in order to have a better tuning.
Resistor CMS 1206 : R108 = 1M (hysteresis resistor).
Resistor 2k added between pin 1 of the J14 connector and the wire (red) going to potentiometer 2k on front panel
Resistor 2k added between pin 3 of the J14 connector and the wire (blue) going to potentiometer 2k on front panel
* Modification of he sample-hold capacitance on the triangular signal:
Capacitor 1206 of 4.7microFarad added on the 1microFarad capacitor (C89).

In the beginning, we measured the optomechanical transfer function of the CC loop with Stanford Research 560 at the beginning. However, since this lock is not so stable, the measurement was very noisy.
Then Pierre put an integrator before 10kHz. We could lock loop with high gain. However, this is a small stable region. If we increase or decrease gain, we can both have oscillation. Anyway, we can lock in that small region. Then we measured both open loop transfer function and optomechanical transfer function.
The result is shown in the attached figure. There are several peaks in the low-frequency region.

Pierre and Yuhang
Here is the open loop transfer function of OPO.

After connecting the power of homodyne and spectrum analyzer(this can remove the ground loop noise), I measured noise spectrum of homodyne again. During this measurement, there is only an infrared beam incident and has a power value of 1.3mW.
The result is shown in the attached figure.

I did a new simulation extending the optical path (of 45 cm), using 4 additional mirrors (fig 1).
The best simulation is shown in fig 2 where the first lens is 47.5 cm far from PBS and the second one is 86.6 cm far from PBS, both with the same focal of 1020 mm.
This solution is more robust than the previous one (entry #1283) as shown in fig 3 and fig 4. Moving the lenses of +/- 1 cm we have a mismatch < 10%.
Next step:
- Simulation for the FC reflected beam (first idea in fig 5).

Yuhang and Matteo
After Pierre replaced the chip of coherent control PLL, we measured the correction signal within a long period of time. This correction signal tells us the stability of CC PLL loop.
The coefficient of variation of the fast correction signal is 0.0538.
While the coefficient of variation of the slow correction signal is 0.1311.
Although we saw a clear change of fast correction signal. The slow correction signal shows much more high-frequency noise.
I also calculated the correlation coefficient between these two signals. It is -0.5155 as we can see one signal is going down while the other is going up.

Since we have new IRMC servo working. I measured again the noise spectrum of homodyne when there is only LO incidence. I did the measurement after common mode noise rejection. The result is shown in the attached figure. There is three main difference.
1. New servo has more flat low-frequency noise.
2. New servo shows much more peaks in the low-frequency region. It looks like electronic noise. As pointed by Matteo, this is because of the ground loop.
3. The peaks at the high frequency of these two situations are different.
After connecting the power of homodyne and spectrum analyzer(this can remove the ground loop noise), I measured noise spectrum of homodyne again. During this measurement, there is only an infrared beam incident and has a power value of 1.3mW.
The result is shown in the attached figure.

[Eleonora P, Yuhang]
Considering the beam parameters of entry #1133, we made the first simulation (using Jammt) for the IR injection telescope. In this simulation (fig 1) we put only two lenses in the straight path on the bench before the FC.
We set as initial position the one of PBS.
Initial beam parameters (entry #1280):
z0 = 15 cm
w0 = 126 um
Resulting beam parameters (entry #1133):
z0 = 4,2673 m
w0 = 1032,7 um
The problem is that this solution is not robust, so there is a large mismatch as function of the position of the lenses of the telescope.
This is shown in fig 2 moving of +/- 5mm the 1st lens and in fig3 moving of +/- 5mm the 2nd lens.
Another problem is that with this configuration we are using the same two mirrors for the FC and homodyne alignment.
Next step:
- extend the optical path before FC using 4 mirrors: good for alignment, bad for losses
- do again the simulation